Liquid water and liquid sulfuric acid mix together to produce aqueous solution of sulfuric acid at 45C. Initially H2O(l) is at 25C temperature and 11kmol/h. H2SO4(l) is at 35C and 1.22kmol/h. Concentration of H2SO4(aq) is 35.7 wt%. Using concentration enthalpy chart, find the total heat needed to transferred during mixing operation.




1
Expert's answer
2021-01-04T03:50:02-0500

M(H2O) = 18 g/mol

M(H2SO4) = 98 g/mol

Concentration enthalpy chart:


"\Delta H_{mix}" (45oC, 35.7%) = -43 kcal/(kg solution) = -180 kJ/(kg solution)

Cp(H2O) = 4.19 kJ/(kg/K)

"\Delta H"(H2O heating) = 4.19*(45-25) = 83.8 kJ/kg

Cp(H2SO4) = 1.34 kJ/(kg/K)

"\Delta H"(H2SO4 heating) = 1.34*(35-25) = 13.4 kJ/kg

Q - total heat needed to transferred (per 1 hour)

Q = -("\Delta H_{mix}" + "\Delta H"(H2O heating) + "\Delta H"(H2SO4 heating)) =

-((11*18+1.22*98)*(-180)+11*18*83.8+1.22*98*13.4) = 38966 kJ/hour